OpenStudy (anonymous):

One of the factors of this trinomial 3x^2 + 12x-36 is A. x + 2 B. x + 6 C. 6 D. x – 3 I think its x+6...am I right?

OpenStudy (turingtest):

yes you are!

OpenStudy (turingtest):

3x^2 + 12x-36 =3(x^2+4x-12)=3(x+6)(x-2)